# Low-Acceleration Gravitational Anomaly from Bayesian 3D Modeling of Wide Binary Orbits: Methodology and Results with Gaia DR3

### v3

@@ 4.3 Future direction of wide binary gravity research
Prior to this work, wide binary gravity research has been based only on observed sky-projected velocity $v_{p}$, and all the previous methods are purely statistical, meaning that an individual binary (except for the specific study of Alpha Centauri: see Banik & Zhao 2018; Banik & Kroupa 2019) does not provide any (even weak) constraint on gravity. Each method considers/proposes a gravity-sensitive quantity $Y$, and derives and compares the observational estimates and theoretical predictions of the medians or distributions of $Y$ in bins of an independent variable $X$, which is usually a proxy of Newtonian acceleration $g_{\rm{N}}(\equiv G_{\rm{N}}M_{\rm{tot}}/r^{2})$ rather than itself. The statistical methods from the literature are summarized in Table 3.
@@
The statistical methods listed in Table 3 have certain limitations. Among others, the actual physical velocity $\mathbf{v}$ is not used and the individual binary-specific eccentricity is usually ignored. For the latter matter, only a series of studies (Chae, 2023a, 2024a, 2024c) by the same author used individual constraints on eccentricities from the Bayesian analysis (Hwang et al., 2022) of the angle between the projected relative displacement and the projected relative velocity. However, these studies are still not fully satisfactory because eccentricities are not self-determined for the gravity tested. These limitations are overcome by the Bayesian 3D modeling introduced in this study. Moreover, the Bayesian 3D modeling naturally takes into account all observational uncertainties including the individually allowed mass uncertainties in inferring gravity.
@@ 5 Conclusion and follow-up work
This work has developed a new Bayesian 3D modeling method for wide binary internal dynamics in the context of testing gravity in low acceleration $\lesssim 10^{-9}$ m s-2. It is shown with mock snapshot observations of wide binary internal dynamics that the method can individually infer probability density functions (PDFs) of the model parameters that correctly encompass the input values of the gravity anomaly parameter $\Gamma$ (Equation (1)) and the orbital parameters. Each PDF of $\Gamma$ is broad given the snapshot nature of the data, but PDFs of $\Gamma$ from a number of independent wide binaries in a similar acceleration (i.e. internal gravity) regime can be combined to result in a much narrower consolidated distribution. The consolidated distribution can be well approximated by the Gaussian function and accurately reproduces the input value of $\Gamma$.
@@
    With 125 wide binaries in a high acceleration range (“Newtonian regime”) $0.080.02<s/r_{\rm{M}}<$ 0.08 ($10^{-7.9}\lesssim g_{\rm{N}}\lesssim 10^{-6.7}$ m s-2), the Bayesian inference returns $\Gamma=0.000\pm 0.011$ or $\gamma_{g}=1.00\pm 0.05$. Thus, internal dynamics of binaries is consistent with Newtonian gravity in acceleration regimes $\gtrsim 100a_{0}$ as far as the first-tier orbital motions are concerned. No first-tier gravitational anomaly is observed in this regime.
@@
    With 76 wide binaries in an intermediate acceleration range (“transition regime”) $0.500.15<s/r_{\rm{M}}<$ 0.50 ($10^{-9.5}\lesssim g_{\rm{N}}\lesssim 10^{-8.5}$ m s-2), the Bayesian inference returns $\Gamma=0.057\pm 0.017$ or $\gamma_{g}=1.30_{-0.09}^{+0.11}$. Newtonian gravity is in tension with these wide binaries at a significance of $3.4\sigma$.
@@
    With 35 wide binaries in a low acceleration range (“MOND regime”) $2.860.50<s/r_{\rm{M}}<$ 2.86 ($10^{-11.0}\lesssim g_{\rm{N}}\lesssim 10^{-9.5}$ m s-2), the Bayesian inference returns $\Gamma=0.085\pm 0.040$ or $\gamma_{g}=1.48_{-0.23}^{+0.30}$. Newtonian gravity is in tension with these wide binaries at a significance of $2.1\sigma$. While the value of $\Gamma$ appears to be higher than in the transition regime, the statistical uncertainty is much larger.
@@
    The inferred values of $\Gamma$ in the MOND and transition+MOND regimes are consistent with a generic value $\Gamma\approx 0.07$ or $\gamma_{g}\approx 1.4$ in the low acceleration ($\lesssim a_{0}$) regime predicted by classical theories of MOND gravity (Bekenstein & Milgrom, 1984; Milgrom, 2010). However, to test a specific theory accurately, numerical simulations of wide binary orbits in the theory are required.
@@
    Eccentricities of orbits inferred by the Bayesian 3D approach agree well with those inferred by Hwang et al. (2022) through an independent Bayesian approach that use only projected quantities and thus do not use radial velocities. In addition, the uncertainties of the new measurements are smaller than Hwang et al. (2022) as a consequence of the additional information from radial velocities.
The present results are limited by the relatively large uncertainties ($\sim 280$ m s-1) of $v_{r}$ that are $\approx 10$ times larger than the uncertainties of the sky-projected velocity components. At the time of this writing, for about 80 wide binaries in a low acceleration regime of $s/r_{\rm{M}}>0.4$, RVs of both stars are being measured with uncertainties $\lesssim 50$ m s-1. With these newly measured RVs, the precision of $\Gamma$ in the MOND regime will improve dramatically. In this study, stars brighter than $M_{G}=3.8$ were excluded to obtain a uniform sample of main-sequence stars by precluding possible contamination by bright subgiant or giant stars. However, as long as involved issues (such as higher probabilities of hiding close companions and availability of accurate masses) can be resolved, it is interesting to consider stars brighter than $M_{G}=3.8$ in the future because brighter stars allow easier measurements of RVs.
@@ Acknowledgments
I thank an anonymous referee for useful and constructive suggestions that led to improved and richer analyses and presentations. I have discussed this work with a number of excellent scientists. In particular, during the summer of 2024 I have visited three institutions and had interesting discussions of this work based on preliminary results. Those scientists include Stacy McGaugh and Tobias Mistele at Case Western Reserve University, Arthur Kosowsky, David Turnshek, and Andrew Zentner at the University of Pittsburgh, and Federico Lelli, Laura Magrini, and Luca Pasquini (a visitor there) at Arcetri Astrophysical Observatory. Luca Pasquini is further thanked for helpful conversations on radial velocities of stars. I thank Bill Janesh at CWRU for help with computing during the summer visit. I thank mathematicians Daniela Calvetti and Erkki Somersalo at CWRU for illuminating discussions on statistical consolidation. I thank Xavier Hernandez and Pavel Kroupa for discussions on wide binary gravity tests. I thank Byeong-Cheol Lee, Dongwook Lim, and David Turnshek for collaboration on measuring radial velocities of stars. I thank Yong Tian and Youngsub Yoon at Sejong for assistance and undertaking new statistical analyses of wide binaries. Y. Yoon is further thanked for carrying out independent calculations of the perspective effects based on the equations of Shaya & Olling (2011) with which the calculations of Appendix B agree well. Finally, I thank Moti Milgrom for encouragements and inspirations of gravity research. Initial results of this work were presented by an invited talk given at the Korean Society of High Energy Physics 2024 Fall Meeting (https://indico.kshep.kr/event/839/contributions/4051). This work was supported by the National Research Foundation of Korea (grant No. NRF-2022R1A2C1092306).
@@ Appendix A Relating the orbital plane to the observer’s frame with the Euler angles
Figure 27 shows an $xyz$ frame including an elliptical orbit lying on its $xy$ plane, arbitrarily oriented with respect to the observer’s $x^{\prime}y^{\prime}z^{\prime}$ frame. The two frames can be related by three successive counterclockwise rotations with Euler angles $\alpha$, $\beta$, and $\gamma$, as shown in (a), (b) and (c) of Figure 27. The angles $\alpha$ and $\gamma$ are rotations about the $z$ and $z^{\prime}$ axes, while the angle $\beta$ is a rotation about an intermediate “$x$” axis (the coinciding $\xi$ and $\xi^{\prime}$ axes in the figure). The three transformations from the initial $xyz$ frame, through the intermediate frames $\xi\eta\zeta$ and $\xi^{\prime}\eta^{\prime}\zeta^{\prime}$, to the final $x^{\prime}y^{\prime}z^{\prime}$ frame can be described by the multiplication of three rotation matrices. The product matrix is given (see Equation (4.46) of Goldstein et al. (2002)) by
@@
In simulating mock snapshot observations of elliptical orbits in the $x^{\prime}y^{\prime}z^{\prime}$ frame, it is sufficient to fix $\alpha=0$ if the periastron is allowed to be arbitrary (i.e. $\phi_{0}$ is random in Figure 1) as is the case in this work. In this case, let $\beta=i$ and $\gamma=\theta$, then the transformation matrix takes a simpler form as
@@
For $\mathbf{\Lambda}=\mathbf{\Lambda}(i)$, Equations (22) and (24) are obtained. Note that $z^{\prime}$, $v_{z^{\prime}}$, $r_{p}(\equiv\sqrt{(x^{\prime})^{2}+(y^{\prime})^{2}}=r\sqrt{\cos^{2}\phi+\sin^{2}\phi\cos^{2}i})$, and $v_{p}(\equiv\sqrt{(v_{x^{\prime}})^{2}+(v_{y^{\prime}})^{2}}=v\sqrt{\cos^{2}\psi+\sin^{2}\psi\cos^{2}i})$ in Equations (A5) and (A6) do not depend on $\theta$ and are the same as those in Equations (22) and (24). This means that only if $r_{p}$ and $v_{p}$ (but not $x^{\prime}$ or $y^{\prime}$ components) are needed, $\mathbf{\Lambda}(i)$ suffices as in previous work (Chae, 2023a, 2024a, 2024c).
@@ Appendix B Perspective effects in proper motions and radial velocities
For a binary system, perspective effects (Shaya & Olling, 2011) refer to apparent relative motions between two stars, observed as biased PMs and RVs, due to the motion of the barycenter with respect to the Sun in the physical 3D space. Perspective effects are expected to be very small for wide binaries in the solar neighborhoods with projected separation $s<20$ kau (El-Badry, 2019). Nevertheless, here I describe a direct numerical algorithm and use it to investigate the perspective effects.
A binary is described using the notation introduced in Section 2.1. With R.A.’s and declinations of two stars $A$ and $B$, the barycenter’s azimuthal and polar angles $\theta$ and $\phi$ (which should not be confused with those used in Appendix A) in the equatorial frame are given by
@@
where all angles are expressed in radians and I assume that the barycenter is the middle between the two stars for simplicity. The system velocity components in the R.A., delc., and radial directions are given by
@@
Now I consider mock physical motions of two stars in the 3D space with the system velocity given by Equation (B3) for time $\Delta t$ into the future (or past). The current positions of the stars $A$ and $B$ are identified by subscript 0 while the future positions by 1. At the current time $t_{0}$, the azimuthal and polar coordinates are obtained by the relations $\phi=\alpha$ and $\theta=\pi/2-\delta$ (with the required conversions of units). The positions of the two stars at $t_{0}$ and future time $t_{1}=t_{0}+\Delta t$ are given in the Cartesian coordinates by
@@
Perspective effects can be a concern only for wide binaries of relatively large separations. Figure 28 shows distributions of perspective effects for 82 wide binaries in a MOND regime ($s/r_{\rm{M}}>0.5$) from the statistical sample (see Table 1). Here perspective effects of PMs are re-expressed as velocities. Because the system velocities (Equation (B2)) are in random directions, the median perspective effects are zero, and only the dispersions are relevant here. For the case $d_{B}=d_{A}$, the standard deviation of all perspective velocities is $37.2$ m s-1, which is comparable to or much smaller than the Gaia DR3 measurement uncertainties of the velocities in the R.A., decl., and radial directions. The perspective velocities normalized by $v$ (which is the magnitude of the observed physical velocity) have a standard deviation of $0.13$. This does not necessarily mean that $v$ is systematically boosted by an appreciable amount because perspective effects can occur in the positive or negative direction for each velocity component. A proper estimate of any systematic bias in the inferred gravity that perspective effects may cause requires individual Bayesian modeling with the PMs and RVs corrected for the perspective effects. Comparison of the second row in Figure 26 with the first row in Figure 22 shows that the effect is very minor (with $\Delta\Gamma=-0.006$ in the MOND regime).
For the case $d_{B}=d_{A}+\mathcal{N}(0,2s)$, the perspective effects have a standard deviation of $\approx 50$ m s-1, which is still comparable to or much smaller than the Gaia DR3 measurement uncertainties of the velocities in the R.A., decl., and radial directions. Moreover, because perspective effects can increase or decrease depending on the distance deviates in individual binaries, the large scatter does not necessarily lead to more reduction of gravitational anomaly than in the case $d_{B}=d_{A}$. To see the effects of the distance deviate $\mathcal{N}(0,2s)$ on $\Gamma$, I have carried out Bayesian modeling of a statistical sample generated by correcting for the perspective effects in PMs and RVs with $d_{B}=d_{A}+\mathcal{N}(0,2s)$. Perhaps ironically, the magnitude of $\Delta\Gamma$ in the MOND regime can be less than in the case $d_{B}=d_{A}$. When the distance difference between the two stars is not measurable, it is perhaps most reliable to assume that it is zero as done in the main body of this paper.

### v4

@@ 4.3 Future direction of wide binary gravity research
Prior to this work, wide binary gravity research has been based only on the observed sky-projected velocity $v_{p}$, and all the previous methods are purely statistical, meaning that an individual binary (except for the specific study of $\alpha$ Centauri; see Banik & Zhao 2018; Banik & Kroupa 2019) does not provide any (even weak) constraint on gravity. Each method considers/proposes a gravity-sensitive quantity $Y$, and derives and compares the observational estimates and theoretical predictions of the medians or distributions of $Y$ in bins of an independent variable $X$, which is usually a proxy of Newtonian acceleration $g_{\rm{N}}(\equiv G_{\rm{N}}M_{\rm{tot}}/r^{2})$ rather than itself. The statistical methods from the literature are summarized in Table 3.
@@
The statistical methods listed in Table 3 have certain limitations. Among others, the actual physical velocity $\mathbf{v}$ is not used and the individual binary-specific eccentricity is usually ignored. For the latter matter, only a series of studies (Chae, 2023a, 2024a, 2024c) by the same author used individual constraints on eccentricities from a Bayesian analysis (Hwang et al., 2022) of the angle between the projected relative displacement and the projected relative velocity. However, these studies are still not fully satisfactory because eccentricities are not self-determined for the gravity tested. These limitations are overcome by the Bayesian 3D modeling introduced in this study. Moreover, the Bayesian 3D modeling naturally takes into account all observational uncertainties including the individually allowed mass uncertainties in inferring gravity.
@@ 5 Conclusion and follow-up work
This work has developed a new Bayesian 3D modeling method for wide binary internal dynamics in the context of testing gravity in low acceleration $\lesssim 10^{-9}$ m s-2. It is shown with mock snapshot observations of wide binary internal dynamics that the method can individually infer PDFs of the model parameters that correctly encompass the input values of the gravity anomaly parameter $\Gamma$ (Equation (1)) and the orbital parameters. Each PDF of $\Gamma$ is broad given the snapshot nature of the data, but PDFs of $\Gamma$ from a number of independent wide binaries in a similar acceleration (i.e., internal gravity) regime can be combined to result in a much narrower consolidated distribution. The consolidated distribution can be well approximated by the Gaussian function and accurately reproduces the input value of $\Gamma$.
@@
    With 125 wide binaries in a high-acceleration range (“Newtonian regime”) $0.080.02<s/r_{\rm{M}}<$ 0.08 ($10^{-7.9}\lesssim g_{\rm{N}}\lesssim 10^{-6.7}$ m s-2), the Bayesian inference returns $\Gamma=0.000\pm 0.011$ or $\gamma_{g}=1.00\pm 0.05$. Thus, internal dynamics of binaries is consistent with Newtonian gravity in acceleration regimes $\gtrsim 100a_{0}$ as far as the first-tier orbital motions are concerned. No first-tier gravitational anomaly is observed in this regime.
@@
    With 76 wide binaries in an intermediate-acceleration range (“transition regime”) $0.500.15<s/r_{\rm{M}}<$ 0.50 ($10^{-9.5}\lesssim g_{\rm{N}}\lesssim 10^{-8.5}$ m s-2), the Bayesian inference returns $\Gamma=0.057\pm 0.017$ or $\gamma_{g}=1.30_{-0.09}^{+0.11}$. Newtonian gravity is in tension with these wide binaries at a significance of $3.4\sigma$.
@@
    With 35 wide binaries in a low-acceleration range (“MOND regime”) $2.860.50<s/r_{\rm{M}}<$ 2.86 ($10^{-11.0}\lesssim g_{\rm{N}}\lesssim 10^{-9.5}$ m s-2), the Bayesian inference returns $\Gamma=0.085\pm 0.040$ or $\gamma_{g}=1.48_{-0.23}^{+0.30}$. Newtonian gravity is in tension with these wide binaries at a significance of $2.1\sigma$. While the value of $\Gamma$ appears to be higher than in the transition regime, the statistical uncertainty is much larger.
@@
    The inferred values of $\Gamma$ in the MOND and transition+MOND regimes are consistent with a generic value $\Gamma\approx 0.07$ or $\gamma_{g}\approx 1.4$ in the low-acceleration ($\lesssim a_{0}$) regime predicted by classical theories of MOND gravity (Bekenstein & Milgrom, 1984; Milgrom, 2010). However, to test a specific theory accurately, numerical simulations of wide binary orbits in the theory are required.
@@
    Eccentricities of orbits inferred by the Bayesian 3D approach agree well with those inferred by Hwang et al. (2022) through an independent Bayesian approach that usees only projected quantities and thus does not use RVs. In addition, the uncertainties of the new measurements are smaller than in Hwang et al. (2022) as a consequence of the additional information from RVs.
The present results are limited by the relatively large uncertainties ($\sim 280$ m s-1) of $v_{r}$ that are $\approx 10$ times larger than the uncertainties of the sky-projected velocity components. At the time of this writing, for a significant number of wide binaries in a MOND regime, RVs with uncertainties $\lesssim 50$ m s-1 are being gathered from various sources/observations. With these newly measured RVs, the precision of $\Gamma$ in the MOND regime will improve dramatically. In this study, stars brighter than $M_{G}=3.8$ were excluded to obtain a uniform sample of main-sequence stars by precluding possible contamination by bright subgiant or giant stars. However, as long as involved issues (such as higher probabilities of hiding close companions and availability of accurate masses) can be resolved, it is interesting to consider stars brighter than $M_{G}=3.8$ in the future because brighter stars allow easier measurements of RVs.
@@ Acknowledgments
I thank the anonymous referee for useful and constructive suggestions that led to improved and richer analyses and presentations. I have discussed this work with a number of excellent scientists. In particular, during the summer of 2024 I have visited three institutions and had interesting discussions of this work based on preliminary results. Those scientists include Stacy McGaugh and Tobias Mistele at Case Western Reserve University, Arthur Kosowsky, David Turnshek, and Andrew Zentner at the University of Pittsburgh, and Federico Lelli, Laura Magrini, and Luca Pasquini (a visitor there) at Arcetri Astrophysical Observatory. Luca Pasquini is further thanked for helpful conversations on the RVs of stars. I thank Bill Janesh at CWRU for help with computing during the summer visit. I thank mathematicians Daniela Calvetti and Erkki Somersalo at CWRU for illuminating discussions on statistical consolidation. I thank Xavier Hernandez and Pavel Kroupa for discussions on wide binary gravity tests. I thank Byeong-Cheol Lee, Dongwook Lim, and David Turnshek for collaboration on measuring the RVs of stars. I thank Yong Tian and Youngsub Yoon at Sejong for assistance and undertaking new statistical analyses of wide binaries. Y. Yoon is further thanked for carrying out independent calculations of the perspective effects based on the equations of Shaya & Olling (2011) with which the calculations of Appendix B agree well. Finally, I thank Moti Milgrom for encouragement and inspirations of gravity research. Initial results of this work were presented by an invited talk given at the Korean Society of High Energy Physics 2024 Fall Meeting (https://indico.kshep.kr/event/839/contributions/4051). This work was supported by the National Research Foundation of Korea (grant No. NRF-2022R1A2C1092306).
@@ Appendix A Relating the orbital plane to the observer’s frame with the Euler angles
Figure 27 shows an $xyz$ frame including an elliptical orbit lying on its $xy$ plane, arbitrarily oriented with respect to the observer’s $x^{\prime}y^{\prime}z^{\prime}$ frame. The two frames can be related by three successive counterclockwise rotations with Euler angles $\alpha$, $\beta$, and $\gamma$, as shown in (a), (b), and (c) of Figure 27. The angles $\alpha$ and $\gamma$ are rotations about the $z$ and $z^{\prime}$ axes, while the angle $\beta$ is a rotation about an intermediate “$x$-axis” (the coinciding $\xi$ and $\xi^{\prime}$ axes in the figure). The three transformations from the initial $xyz$ frame, through the intermediate frames $\xi\eta\zeta$ and $\xi^{\prime}\eta^{\prime}\zeta^{\prime}$, to the final $x^{\prime}y^{\prime}z^{\prime}$ frame can be described by the multiplication of three rotation matrices. The product matrix is given (see Equation (4.46) of Goldstein et al. (2002)) by
@@
In simulating mock snapshot observations of elliptical orbits in the $x^{\prime}y^{\prime}z^{\prime}$ frame, it is sufficient to fix $\alpha=0$ if the periastron is allowed to be arbitrary (i.e., $\phi_{0}$ is random in Figure 1) as is the case in this work. In this case, let $\beta=i$ and $\gamma=\theta$, then the transformation matrix takes a simpler form as
@@
For $\mathbf{\Lambda}=\mathbf{\Lambda}(i)$, Equations (22) and (24) are obtained. Note that $z^{\prime}$, $v_{z^{\prime}}$, $r_{p}(\equiv\sqrt{(x^{\prime})^{2}+(y^{\prime})^{2}}=r\sqrt{\cos^{2}\phi+\sin^{2}\phi\cos^{2}i})$, and $v_{p}(\equiv\sqrt{(v_{x^{\prime}})^{2}+(v_{y^{\prime}})^{2}}=v\sqrt{\cos^{2}\psi+\sin^{2}\psi\cos^{2}i})$ in Equations (A5) and (A6) do not depend on $\theta$ and are the same as those in Equations (22) and (24). This means that only if $r_{p}$ and $v_{p}$ (but not $x^{\prime}$ or $y^{\prime}$-components) are needed, $\mathbf{\Lambda}(i)$ suffices as in previous work (Chae, 2023a, 2024a, 2024c).
@@ Appendix B Perspective effects in proper motions and radial velocities
For a binary system, perspective effects (Shaya & Olling, 2011) refer to the apparent relative motions between two stars, observed as biased PMs and RVs, due to the motion of the barycenter with respect to the Sun in the physical 3D space. Perspective effects are expected to be very small for wide binaries in the solar neighborhood with projected separation $s<20$ kau (El-Badry, 2019). Nevertheless, here I describe a direct numerical algorithm and use it to investigate the perspective effects.
A binary is described using the notation introduced in Section 2.1. With the R.A. and decl. of two stars $A$ and $B$, the barycenter’s azimuthal and polar angles $\theta$ and $\phi$ (which should not be confused with those used in Appendix A) in the equatorial frame are given by
@@
where all angles are expressed in radians and I assume that the barycenter is the middle between the two stars for simplicity. The system velocity components in the R.A., decl., and radial directions are given by
@@
Now I consider mock physical motions of two stars in the 3D space with the system velocity given by Equation (B3) for time $\Delta t$ into the future (or past). The current positions of the stars $A$ and $B$ are identified by subscript 0 while the future positions by subscript 1. At the current time $t_{0}$, the azimuthal and polar coordinates are obtained by the relations $\phi=\alpha$ and $\theta=\pi/2-\delta$ (with the required conversions of units). The positions of the two stars at $t_{0}$ and future time $t_{1}=t_{0}+\Delta t$ are given in the Cartesian coordinates by
@@
Perspective effects can be a concern only for wide binaries of relatively large separations. Figure 28 shows distributions of perspective effects for 82 wide binaries in the MOND regime ($s/r_{\rm{M}}>0.5$) from the statistical sample (see Table 1). Here perspective effects of PMs are reexpressed as velocities. Because the system velocities (Equation (B2)) are in random directions, the median perspective effects are zero, and only the dispersions are relevant here. For the case $d_{B}=d_{A}$, the standard deviation of all perspective velocities is $37.2$ m s-1, which is comparable to or much smaller than the Gaia DR3 measurement uncertainties of the velocities in the R.A., decl., and radial directions. The perspective velocities normalized by $v$ (which is the magnitude of the observed physical velocity) have a standard deviation of $0.13$. This does not necessarily mean that $v$ is systematically boosted by an appreciable amount because perspective effects can occur in the positive or negative direction for each velocity component. A proper estimate of any systematic bias in the inferred gravity that perspective effects may cause requires individual Bayesian modeling with the PMs and RVs corrected for the perspective effects. Comparison of the second row in Figure 26 with the first row in Figure 22 shows that the effect is very minor (with $\Delta\Gamma=-0.006$ in the MOND regime).
For the case $d_{B}=d_{A}+\mathcal{N}(0,2s)$, the perspective effects have a standard deviation of $\approx 50$ m s-1, which is still comparable to or much smaller than the Gaia DR3 measurement uncertainties of the velocities in the R.A., decl., and radial directions. Moreover, because perspective effects can increase or decrease depending on the distance deviates in individual binaries, the large scatter does not necessarily lead to a greater reduction of the gravitational anomaly than in the case $d_{B}=d_{A}$. To see the effects of the distance deviate $\mathcal{N}(0,2s)$ on $\Gamma$, I have carried out Bayesian modeling of a statistical sample generated by correcting for the perspective effects in PMs and RVs with $d_{B}=d_{A}+\mathcal{N}(0,2s)$. Perhaps ironically, the magnitude of $\Delta\Gamma$ in the MOND regime can be less than in the case $d_{B}=d_{A}$. When the distance difference between the two stars is not measurable, it is perhaps most reliable to assume that it is zero as done in the main body of this paper.